# Culture is Not Trivia: Sociocultural Theory for Cultural NLP

### v1

@@ Abstract
Naitian Zhou,1 David Bamman1and Isaac L. Bleaman2 1School of Information, University of California, Berkeley 2Department of Linguistics, University of California, Berkeley {naitian,dbamman,bleaman}@berkeley.edu,
@@ 2 The goals of cultural NLP
There has been relatively limited exploration in building systems that are explicitly reflexive Sorensen et al. (2024). Instead, a wide body of work focuses on assessing whether NLP systems can generate different, appropriate outputs in response to different cultural contexts — by answering value-oriented survey questions in a manner consistent with the target culture, for example ( Cao et al., 2024a; Huang and Yang, 2023). Some works that center adaptation as a value focus on extrinsic evaluation: instead of probing whether language models know specific cultural facts, they test whether NLP systems respond in a way that demonstrates this knowledge Bhatt and Diaz (2024).
@@
In addition to breadth, there is a desire for depth in the form of more granular and extensive cultural understanding. The value of nuance motivates works which build resources for specific languages or resources Koto et al. (2024); Son et al. (2024); Li et al. (2024b) and explore locally meaningful cultural categories ( Dev et al., 2023). These works might rely on local informants to provide cultural knowledge Koto et al. (2024), trading breadth of coverage for richer cultural knowledge that large-scale survey-based methods cannot capture.
Though these four desiderata are not mutually exclusive, they coalesce into two sets of values that are largely unrelated (visualized in fig. 1). The first two desiderata reflect two kinds of cultural competence: the knowledge of how to respond differentially, and the knowledge of when it is appropriate to do so. The second two desiderata reflect an orthogonal value of cultural coverage: we want systems that cover many cultures, as well as systems that cover many aspects of each culture.
@@ 3 Recurring troubles
The most commonly cited limitation was one of coverage (40% of papers): the dataset or evaluation being presented was only collected with respect to a small subset of cultures. Partly, this can be explained by the proxies being used for setting cultural boundaries. Of the papers we surveyed, 36% of them used nationality as a demographic proxy Adilazuarda et al. (2024). However, many papers problematize this choice in the limitations section, since nations are politically defined and not culturally homogeneous Bickham et al. , and language labels usually reflect a hegemonic notion of a standard variety Lippi-Green (2011).
Another common limitation was a lack of dynamicity (12% of papers): culture is constantly constructed through social negotiation Ochs (2009), but benchmarks are largely static collections of examples or facts Son et al. (2024); Keleg and Magdy (2023); Jin et al. (2024); Li et al. (2024b). In most works, there is no granularity in the temporal dimension, failing to achieve an aspect of the desired nuance. Uncertainty around the definition of culture also limits nuance in cultural technologies, since most papers focus only on a small subset of culturally-relevant objects through proxies like food, etiquette, or values, without a framework to unify them. Roughly 37% of papers directly problematize their choice of a particular cultural proxy as being limited in its ability to represent culture as a whole, or too coarse to capture intragroup variation (28% of papers mention this specifically).
@@ 4.1 Text and context
We take this framework as the point of departure for the rest of this paper. We describe the foundational concepts (§4.2) and provide a case study for how they can clarify the objects and goals of cultural NLP (§5). Then, we take a broader look at how sociocultural linguistic theory can inform computational work on cultural competence (§6.1) and advance the goals of cultural NLP (§6.2).
@@ 4.2 A primer on sociocultural linguistics
Bucholtz and Hall (2005) lay out five core principles of sociocultural linguistics. The terminology they use centers on the idea of identity as the “social positioning of self and other.” As we explained above, this is a useful way of conceptualizing the cultural system more generally.
@@
Identity includes multiple levels of categories, including macro-level demographics, locally-specific meaning, and contextually-specific stances and styles. This is a more inclusive and nuanced notion of culture that speaks to one of the core limitations of current work. National identity is only one level at which identity occurs, and the idea of positionality insists that we understand more granular categories of identity as well, including ones that are local to a specific community or even an interaction.
@@ 5 Case study: culture from text
In §4.1, we distinguish between language (the text) as the traditional object of study in linguistics and various ways of assessing culture as the surrounding social context. It has become essentially paradigmatic within NLP that we should expect to derive extratextual information (such as world models, for example) from training on text alone. It is reasonable, then, that there is a vein of cultural NLP work that mines for facts about specific identities from culturally-centered discourse on social media or the Internet Shi et al. (2024); Fung et al. (2024); Nguyen et al. (2023). In some cases, large language models are prompted to generate specific culturally-relevant scenarios Qiu et al. (2024b). In many of these papers, authors note the dangerous potential for extracting biased or stereotyping information. How should we understand the epistemic status of these surfaced facts as cultural knowledge? In this section, we apply indexical theory to demonstrate that these works, in fact, can only learn stereotypes.
These papers aim to construct an indexical field. Cultural facts aggregated by these systems generally resemble the form:
@@ 5.1 Stereotypes all the way down
The implication of this idea is that works which study cultural discourse are primarily studying the stereotypes embedded in the ideologies of the groups that generated this data, and only incidentally studying the cultures that are the objects of discourse. In fact, Labov (1973) defines stereotype exactly as the linguistic forms which are subject to metapragmatic discussion.
@@ 6.2 Localized NLP
All told, we are far from building culturally competent systems, given these clear and pressing theoretical and methodological gaps. But cultural NLP also faces more immediate goals, which are perhaps more central to the field as it is currently configured. We want to create, e.g., web agents that will not make food purchases that violate religious dietary laws Qiu et al. (2024a) or image generation models that show the local currency when displaying money Khanuja et al. (2024). Do we need to achieve cultural competence in the general sense for these more immediate applications?
@@
Thus, localization allows us to focus on particulars that may be more tractably implemented and evaluated in real-world systems today.
@@ Appendix A List of surveyed works
Evaluating Cultural and Social Awareness of LLM Web Agents Qiu et al. (2024a) multilingual, coverage, context
@@ Appendix B Category key
based on existing model Notes that automated evaluation is based on existing model.

### v2

@@ Abstract
Naitian Zhou,1 David Bamman1and Isaac L. Bleaman2 1School of Information, University of California, Berkeley 2Department of Linguistics, University of California, Berkeley {naitian,dbamman,bleaman}@berkeley.edu
@@ 2 The goals of cultural NLP
A wide body of work focuses on assessing whether NLP systems can generate different, appropriate outputs in response to different cultural contexts — by answering value-oriented survey questions in a manner consistent with the target culture, for example ( Cao et al., 2024a; Huang and Yang, 2023). Some works that center adaptation as a value focus on extrinsic evaluation: instead of probing whether language models know specific cultural facts, they test whether NLP systems respond in a way that demonstrates this knowledge Bhatt and Diaz (2024).
@@
In addition to breadth, there is a desire for depth in the form of more granular and extensive cultural understanding. The value of nuance motivates works which build resources for specific languages or contexts Koto et al. (2024); Son et al. (2025); Li et al. (2024b) and explore locally meaningful cultural categories ( Dev et al., 2023). These works might rely on local informants to provide cultural knowledge Koto et al. (2024), trading breadth of coverage for richer cultural knowledge that large-scale survey-based methods cannot capture.
Though these four desiderata are not mutually exclusive, they roughly coalesce into two sets of trade-offs (visualized in fig. 1). The first two desiderata reflect two kinds of cultural competence: the knowledge of how to respond differentially, and the knowledge of when it is appropriate to do so. The second two desiderata reflect an orthogonal value of cultural coverage: we want systems that cover many cultures, as well as systems that cover many aspects of each culture.
@@ 3 Recurring troubles
The most commonly cited limitation was one of coverage (40% of papers): the dataset or evaluation being presented was only collected with respect to a small subset of cultures. Partly, this can be explained by the proxies being used for setting cultural boundaries. Of the papers we surveyed, 36% of them used nationality as a demographic proxy Adilazuarda et al. (2024). However, many papers problematize this choice in the limitations section as lacking nuance, since nations are politically defined and not culturally homogeneous Méndez and Naples (2014), and language labels usually reflect a hegemonic notion of a standard variety Lippi-Green (2012).
Another common limitation was a lack of dynamicity (12% of papers): culture is constantly constructed through social negotiation Ochs (2009), but benchmarks are largely static collections of examples or facts Son et al. (2025); Keleg and Magdy (2023); Jin et al. (2024); Li et al. (2024b). In most works, there is no granularity in the temporal dimension, failing to achieve an aspect of the desired nuance. Uncertainty around the definition of culture also limits nuance and inclusivity in cultural technologies, since most papers focus only on a small subset of culturally-relevant objects through proxies like food, etiquette, or values, without a framework to unify them. Roughly 37% of papers directly problematize their choice of a particular cultural proxy as being limited in its ability to represent culture as a whole, or too coarse to capture intragroup variation (28% of papers mention this specifically).
@@ 4.1 Text and context
We take this framework as the point of departure for the rest of this paper. We describe some foundational concepts (§4.2) and provide a case study for how they can clarify the objects and goals of cultural NLP (§5). Then, we take a broader look at how sociocultural linguistic theory can inform computational work on cultural competence (§6.1) and advance the goals of cultural NLP (§6.2).
@@ 4.2 A primer on sociocultural linguistics
Bucholtz and Hall (2005) lay out five core principles of sociocultural linguistics. The terminology they use centers on the concept of identity, defined as “the social positioning of self and other.” If we take the analogy of identity as a social position, we might consider culture to be the broader landscape within which identities are located. Through this lens, it becomes clear that individual identities not only reflect, but also constitute culture, giving shape to the cultural terrain. Thus, sociocultural linguistics provides a useful way of conceptualizing the cultural system more generally. [^4] While Bucholtz and Hall (2005) is nominally centered around identity and interaction, they in fact discuss culture more generally at various points (e.g., when introducing partialness).
@@
Identity includes multiple levels of categories, including macro-level demographics, locally-specific distinctions, and contextually-specific stances and styles. This is a more inclusive and nuanced notion of culture that speaks to one of the core limitations of current work. National identity is only one level at which identity occurs, and the idea of positionality insists that we understand more granular categories of identity as well, including ones that are local to a specific community or even an interaction.
@@ 5 Case study: culture from text
In §4.1, we distinguish between language (the text) as the traditional object of study in linguistics and various ways of assessing culture as the surrounding social context. It has become essentially paradigmatic within NLP that we should expect to derive extratextual information (such as world models, for example) from training on text alone. It is reasonable, then, that there is a vein of cultural NLP work that mines for facts about specific identities from culturally-centered discourse on social media or the Internet Rao et al. (2025); Fung et al. (2024); Nguyen et al. (2023). In some cases, large language models are prompted to generate specific culturally-relevant scenarios Qiu et al. (2024). In many of these papers, authors note the dangerous potential for extracting biased or stereotyping information. How should we understand the epistemic status of these surfaced facts as cultural knowledge? Labov (1972, 314) defines stereotype as the linguistic forms which are subject to metapragmatic discussion: that is, the signs whose meanings are actively discussed. In this section, we apply indexical theory to demonstrate that these works, in fact, can only learn stereotypes.
Papers that mine for cultural facts aim to construct an indexical field from unstructured web text. These systems take as input online documents or discussions about cultural differences (e.g., a guide to dining etiquette in India; Rao et al., 2025) and extract indexical mappings resembling this form:
@@ 5.1 Stereotypes all the way down
The implication of this idea is that works which study cultural discourse are primarily studying the stereotypes embedded in the ideologies of the groups that generated this data, and only incidentally studying the cultures that are the objects of discourse. The status of these cultural facts as stereotypes does not depend on whether the group generating the discourse is the same as the group serving as the subject: the fact that these are cultural associations being discussed rather than observed classifies them as stereotypes. In fact, groups can, and often do, have both positive and negative stereotypes for themselves Leavitt et al. (2015); Coffman (2014); Sinclair et al. (2006); Pickett et al. (2002).
@@
This is also not to say that first-order indices are inaccessible through computational methods. For example, computational sociolinguistic work successfully identifies first-order social meaning by using platform metadata like Twitter geolocation Grieve et al. (2019) or subreddit Zhang et al. (2017); Lucy and Bamman (2021) to measure sociolinguistic variation in linguistic features (often lexical or semantic) and the sociocultural meaning they index (like locale or community).
@@ 6.2 Localized NLP
All told, we are far from building culturally competent systems, given these clear and pressing theoretical and methodological gaps. But cultural NLP also faces more immediate goals, which are perhaps more central to the field as it is currently configured. We want to create, e.g., web agents that will not make food purchases that violate religious dietary laws Qiu et al. (2025) or image generation models that show the local currency when displaying money Khanuja et al. (2024). These goals are mediated by constraints on resources: user studies and qualitative interviews might provide richer cultural data, but are more expensive to scale. Do we need to achieve cultural competence in the general sense for these more immediate applications? And how do we navigate trade-offs between desiderata?
@@
Thus, localization allows us to focus on particulars that may be more tractably implemented and evaluated in real-world systems today. Principles from user-centered Lowdermilk (2013) and participatory design Caselli et al. (2021), particularly as applied to machine learning contexts Ayobi et al. (2023); Tan et al. (2024), provide footholds for NLP practitioners who wish to identify the cultural features which are relevant to a given application context.
@@ 8 Limitations
Acknowledgments
This work was supported by funding from the National Science Foundation (Graduate Research Fellowship DGE-2146752).
We thank Laura Sterponi, Kent Chang, Dan Hickey, Nikita Mehandru, Lucy Li, and many others for their thoughtful discussion and feedback.
@@ Appendix A List of surveyed works
Evaluating Cultural and Social Awareness of LLM Web Agents Qiu et al. (2025) multilingual, coverage, context
@@ Appendix B Category key
based on existing model Notes that automated evaluation is based on an existing model.